OpenStudy (anonymous):

Suppose that f and g are continuous on [a,b] and differentiable on (a,b) and suppose that f(a)=g(a) and f'(x) < g'(x) for a

OpenStudy (paxpolaris):

are you familiar with mean value theorem ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

if f is continuous on the closed interval [a,b] and differentiable on the open interval (a,b) there there is at least one number c in (a,b) such that f'(c) = f(b)-f(a)/b-a

OpenStudy (paxpolaris):

nvm .. you can't actually use it.

OpenStudy (anonymous):

actually i think you can?

OpenStudy (anonymous):

wait no... i dont know what im talking about

OpenStudy (paxpolaris):

how about: \[f'(x)<g'(x)\] can we take the definite integral from a to b....

OpenStudy (paxpolaris):

can we do this:\[\int\limits_a^b f'(x) < \int\limits_a^b g'(x)\]

OpenStudy (anonymous):

thanks for your help

OpenStudy (anonymous):

I don't think integrals are necessary here. You're explicitly told that \(f'(x)<g'(x)\), so \(f\) is decreasing faster (if both derivatives are negative) or increasing slower (if both are positive) than \(g\), and so \(f(b)<g(b)\) provided that \(a<b\).

OpenStudy (anonymous):

One more case to consider, actually, however trivial it may seem. If \(f'<0\) and \(g'>0\), then \(g\) is increasing faster than \(f\), so again \(f(b)<g(b)\).